OpenStudy (anonymous):

I don't understand this question, please help: Write an equation of the circle whose diameter AB has endpoints A(-4,2) and B(4,-4).

OpenStudy (anonymous):

find the midpoint of the line segment, that will be the center do you know how to find the midpoint?

OpenStudy (anonymous):

I think so, use the midpoint formula right?

OpenStudy (anonymous):

right, take the average in each coordinate (i.e. add up and divide by 2) in this case you can just about do it in your head

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

Okay I will, hang on.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I got my H and K value to be (0,-1)

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

now we know it is going to look like \[(x-0)^2+(y-(-1))^2=r^2\] or \[x^2+(y+1)^2=r^2\] so we now need \(r\) right?

OpenStudy (anonymous):

and since we have the midpoint and also a point on the circle (actually two points) we can use the distance formula to find \(r^2\) you know the distance formula?

OpenStudy (anonymous):

yes i know that formula, hang on so i can write this down

OpenStudy (anonymous):

k

OpenStudy (anonymous):

for the distance i got d to equal 100

OpenStudy (anonymous):

seems rather unlikely lets check

OpenStudy (anonymous):

you want the distance between the two points \((0,-1)\) and \((4,-4)\) right?

OpenStudy (anonymous):

oh, i messed up. the distance should have equaled 65? if I'm finding the distance between (0, -1) and (4, -4)...

OpenStudy (anonymous):

yes, actually you do not need the distance, just the square of the distance because you need \(r^2\) not \(r\) but that is not right either we can use the formula, but first we can think

OpenStudy (anonymous):

from 0 to 4 is 4 units in the \(x\) direction, and from -1 to -4 is 3 units in the \(y\) direction, and so by pythagoras the square of the distance is \(3^2+4^2=9+16=25\)

OpenStudy (anonymous):

here we are just using the old formula for a right triangle \(a^2+b^2=c^2\) which is the basis for the distance formula

OpenStudy (anonymous):

if you want to use the distance formula it is \[\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] in this case you get \[\sqrt{4-0)^2+(-1+4)^2}=\sqrt{4^2+3^2}=\sqrt{25}=5\]

OpenStudy (anonymous):

i am not sure how you arrived at 65, but i hope it is clear what the answer is \[r=5, r^2=25\]

OpenStudy (anonymous):

and your "final answer" is therefore \[x^2+(y+1)^2=25\] that is the equation for this circle